# Crowd-FM: Learned Optimal Selection of Conditional Flow Matching-generated Trajectories for Crowd Navigation

### v1

@@ IV-C Real-World Experiments
We validated Crowd-FM in real-world scenarios on two distinct robots: a Husky A200 and a custom-made autonomous wheelchair. Each robot was equipped with a Livox Mid-360 3D LiDAR, from which we extracted 2D planar point clouds for our processing pipeline and an Intel® Realsense D455 RGBD Camera to estimate dynamic obstacle velocities and trajectories using the LV-DOT algorithm [14]. The experiments were conducted indoors and outdoors, where our method demonstrated consistently reasonable success rates. We demonstrate some testing results in Fig 7.
Figure 7: Crowd-FM in action in the real world, both on a custom wheelchair prototype and a Clearpath Husky A200.
@@ V CONCLUSION AND FUTURE WORK
This work introduces Crowd-FM, a trajectory planning framework that utilises Conditional Flow Matching to learn expressive distributions over Bernstein polynomial control points. Unlike prior generative approaches, which have the drawback of discrete latent bottlenecks (VQ-VAEs) or are computationally expensive (Diffusion), Crowd-FM demonstrates that flow-based transport can capture diverse and smooth navigation, while remaining efficient at inference. A lightweight optimizer further refines the candidate trajectories, boosting success by $\sim$20%, without resorting to a slower complete trajectory optimization. We validated the approach on both a simulated hand-made dataset, as well as real-world testing on two separate robots, highlighting that smooth, non-freezing trajectories can be effectively reproduced and generalised in dense crowd scenarios.
We also see potential in coupling Flow-based trajectory generation with complementary planning modules, such as predictive human motion models or model predictive controllers, to study how we can leverage the strengths of both methods. Another important direction is to expand the dataset with more diverse and heterogeneous scenarios, letting us capture richer behaviours and further test the scoring function under different crowd dynamics.

### v2

@@ IV-C Real-World Experiments
We validated Crowd-FM in real-world scenarios on a Pioneer 3-DX mobile robot. The robot was equipped with an RPLiDAR 2D scanner and utilized a LegTracker[8] module to estimate dynamic obstacle velocities and positions. The experiments were conducted in our lab and in the academic block, where our method consistently demonstrated reasonable success rates in densely populated environments. In these complex settings, the system achieved an overall success rate of 17 runs out of 20 total trials. Fig. 7 shows stills from two such representative testing scenarios.
Figure 7: Crowd-FM in action on a P3-DX robot in two distinct scenarios - a cluttered environment with high static obstacle density(lab), and a crowded environment with high dynamic obstacle density(academic block).
@@ V CONCLUSION AND FUTURE WORK
This work introduces Crowd-FM, a trajectory-planning framework that utilizes Conditional Flow Matching to learn expressive distributions over Bernstein polynomial control points. Unlike prior generative approaches, which have discrete latent bottlenecks(VQ-VAEs) or are computationally expensive(Diffusion), Crowd-FM demonstrates that flow-based transport can capture diverse and smooth navigation, while remaining efficient at inference. A lightweight projection optimization step further refines the candidate trajectories, boosting success by $\sim$20%, without resorting to a slower complete trajectory optimization. We validate the approach both in simulation and real-world settings, highlighting that smooth, non-freezing, and human-like trajectories can be rapidly generated in dense crowds.
An essential future direction is to expand the dataset with more heterogeneous scenarios, allowing us to capture richer behaviors and test the scoring function under different crowd dynamics. We also see CFM’s potential in learning social cues and scene semantics through its integration with complementary planning modules, such as human motion prediction models, moving towards social navigation.
VI Acknowledgment
The authors would like to acknowledge the use of large language models (Gemini, ChatGPT) and writing enhancement tools (Grammarly) solely for refining the grammatical structure and the linguistic flow of this manuscript.